# Complete NLO BFKL impact factors for quarkonium hadroproduction in NRQCD: the case of $^{1}S_{0}^{[1]}$, $^{1}S_{0}^{[8]}$, and $^{3}S_{1}^{[8]}$ states

### v1

@@ Abstract
We present the first complete next-to-leading order calculation of the impact factors for hadroproduction of $S$-wave quarkonium states within the BFKL formalism. We present the computation of the real-emission contributions which completes the recent one of one-loop virtual corrections by one of us for the impact factors for the $^{1}S_{0}^{[1]}$, $^{1}S_{0}^{[8]}$, and $^{3}S_{1}^{[8]}$ NRQCD states. We prove the cancellation of soft divergences between real and virtual contributions, and that the surviving collinear singularities are compatible with factorisation up to one loop for a novel class of processes where BFKL resummation can be applied. Our work indeed represents the first complete NLO quarkonium impact factor in the BFKL framework and paves the way to first next-to-leading-logarithmic-precision studies for hadroproduction of forward-backward quarkonium associated production at hadron colliders.
@@ 5.1 Extracting divergences from the CS subtraction term
$V_{q\text{, ST-2}}^{(m,1)}=C_{\overline{MS}}z|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})\bigg(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\bigg)^{\epsilon}\bigg(-\frac{1}{\epsilon}+O(\epsilon)\bigg)\left[P_{gq}(z)-\epsilon C_{F}z\right],$ (51)
@@ A.1 LLA and Multi-Regge Kinematics
with $q_{1,2}^{\mu}=x_{1,2}P_{1,2}^{\mu}$. In the LLA case the initial partons $a=b=g$ and final-state partons are in Multi-Regge Kinematics (MRK):
@@
which is the kinematic region which generates the leading logarithmic contribution to the cross section. We start with the standard expression for the partonic cross section in QCD:
@@
where the ${\cal A}_{\text{MRK}}$ is the ordinary tree-level QCD amplitude of the process (106) in the MRK limit (107)+(108), the factors $N_{a}^{\text{(pol.)}}$ stand for the averaging over helicities of the incoming on-shell parton ($N_{g}^{\text{(pol.)}}=D-2=2-2\epsilon$ and $N_{q}^{\text{(pol.)}}=2$), the $N_{a}^{\text{(col.)}}$ average over colours of the parton ($N_{q}^{\text{(col.)}}=N_{c}$ and $N_{g}^{\text{(col.)}}=N_{c}^{2}-1$), the partonic flux-factor is $I_{ab}=2(q_{1}+q_{2})^{2}=2q_{1}^{+}q_{2}^{-}$ and $\delta_{+}(p^{2}-m^{2})=\delta(p^{2}-m^{2})\theta(p^{0})$.
Due to conditions (107) and (108) the momentum-conservation $\delta$-function can be approximated as:
@@
For the final-state particles one transitions to the transverse momentum and rapidity variables in a standard way, e.g.:
@@
The crucial step of course is the factorisation for ${\cal A}_{\text{MRK}}$ in the kinematic regime (112), which is illustrated diagrammatically in the Fig. 3(a). The corresponding factorised form holds up to corrections $O(e^{y_{k}-y_{1}})$ and $O(e^{y_{2}-y_{k}})$ and can be written in terms of reggeised gluon exchanges and vertices with the help of gauge-invariant EFT for multi-Regge processes in QCD Lipatov95 :
@@
Figure 3: Diagrammatic representation for the factorisation of the amplitude (106) in the MRK (113) and QMRK (133). The dashed lines denote reggeised gluon propagators.
@@
Thus for the squared amplitude (113) one obtains:
@@
Inserting this result into eqn. (109), introducing the variables $z_{1}=p_{1}^{+}/q_{1}^{+}$, $z_{2}=p_{2}^{-}/q_{2}^{-}$ and taking into account approximations (110) one can rewrite the partonic cross section in a form of eqn. (7) [^11] Note the difference of measures $d^{2}{\bm{p}}_{T}$ vs. $d|{\bm{p}}_{T}|d\phi$ w.r.t. eqn. (7).:
@@
Where, taking into account the NRQCD-factorisation, the LO IFs are defined as:
@@
Also in the eqn. (118) the $O(\alpha_{s})$ real-emission contribution to the BFKL Green’s function appears:
@@
which can be understood as a single iteration of the real-emission part of the LO BFKL kernel, see the eqn. (122) below.
@@ A.2 Different forms of the LO BFKL kernel
To see the cancellation of IR-divergence between real-emission and virtual terms of eqn. (122) let’s convolute it with a smooth test-function $G({\bm{q}}_{T})$ and do a few transformations:
@@
In order for the integral in the second line of eqn. (124) to be explicitly IR and UV finite, the function $\Theta({\bm{p}}_{T},{\bm{q}}_{T})$ has to satisfy two conditions: (i) it should be equal to $1$ for ${\bm{p}}_{T}={\bm{q}}_{T}$ in order for IR-divergence to cancel, (ii) it should decrease at ${\bm{q}}_{T}^{2}\gg{\bm{p}}_{T}^{2}$ in order to avoid appearance of an artificial UV-divergence at large ${\bm{q}}_{T}$. Of course there exist infinitely many such functions, we will consider the following several examples:
@@
$\Theta_{3}({\bm{p}}_{T},{\bm{q}}_{T})=\frac{{\bm{p}}_{T}^{2}}{{\bm{q}}_{T}^{2}+({\bm{p}}_{T}-{\bm{q}}_{T})^{2}},$ (127)
Let’s compute the integrals corresponding to the $\Theta_{1}$ and $\Theta_{2}$:
@@
where we have put ${\bm{q}}_{T}={\bm{p}}_{T}+{\bm{k}}_{T}$. So one can see that for both of these choices the integral of $\Theta$ in the last line of eqn. (124) cancels the virtual contribution up to $O(\epsilon)$ and for $\epsilon=0$ only the second line of eqn. (124) survives. Therefore for the case of $\Theta_{1}$ one recovers the eqn. (26) and it corresponds to our default representation of the LO BFKL kernel (25).
@@
So in the end of the day, the integral of $\Theta_{3}$ also cancels the virtual part of the kernel up to $O(\epsilon)$ the corresponding explicitly IR-finite form of the kernel is used e.g. in Forshaw:1997dc (eqn. (4.18), Sec. 4.3).
@@ A.3 NLLA and Quasi-Multi-Regge Kinematics
One of the contributions to the NLLA is when the MRK conditions (107) and (108) are violated and the emitted gluon carries-away a significant fraction of momentum of the incoming parton $a$ in the process (106):
@@
This is one of the examples of Quasi-MRK (QMRK). In this kinematic limit the squared tree-level QCD amplitude of the process (106) factorises as:
@@
with the Reggeon transverse-momentum ${\bm{q}}_{T}=-{\bm{p}}_{2T}={\bm{p}}_{1T}+{\bm{k}}_{T}$, see the Fig. 3(b). The corresponding approximation for momentum-conservation $\delta$-function is:
@@
since we can no longer neglect the gluon’s $(+)$ momentum component. Substituting approximations (133) and (134) to the eqn. (109) and rewriting it in a form of eqn. (7) one obtains:
@@
where the integrals over ${\bm{k}}_{T}$ and $k^{+}$ can be calculated with the help of $\delta$-functions, leaving behind a rational function of ${\bm{q}}_{T}$, ${\bm{p}}_{T}$ and $z$.
For convenience, we express the NLO real-emission correction by factoring-out the LO IF $h_{m}^{(0)}({\bm{p}}_{T})$ and the factor of $\alpha_{s}/(2\pi)$ (see eqns. (10) and (17)):
@@ Appendix B Exact real-emission squared matrix elements
In this appendix we will collect the results for exact squared matrix elements of the processes:
@@
where $R_{-}$ is the reggeised gluon carrying momentum $q^{\mu}_{R}=n_{+}^{\mu}q_{R}^{-}/2+q_{T}^{\mu}$, which is off-shell: $q_{R}^{2}=-{\bm{q}}_{T}^{2}$, the parton $a=g/q$ is on-shell ($q^{\mu}=q^{+}n_{-}^{\mu}/2$, $q^{2}=0$) and the NRQCD states of the $Q\bar{Q}$-pair carrying the momentum $p$ ($p^{2}=M^{2}$) are $m={}^{1}S_{0}^{[1]}$, $^{1}S_{0}^{[8]}$ and $^{3}S_{1}^{[8]}$. The amplitude of the process (140) is computed according to the Feynman rules of the gauge-invariant EFT for Multi-Regge processes in QCD Lipatov95 . The relevant Reggeon-gluon couplings are summarised in the Sec. 4 of ref. Nefedov:2024swu .
The projection of the $Q\bar{Q}$-pair on the corresponding NRQCD intermediate state is done in a standard way. To project-out the $L=0$-state we simply put $p_{Q}=p_{\bar{Q}}=p/2$. The heavy quark spinors are substituted by the corresponding spin and colour projectors as:
@@
with $T^{a}_{ij}$ being the $SU(N_{c})$-generators in the fundamental representation. To obtain the hard-scattering coefficient in accordance with the eqn. (9) one has to multiply the projected squared amplitude by $M^{3}/(N_{\text{col.}}^{\text{(NRQCD)}}N_{\text{pol.}}^{\text{(NRQCD)}})$ with $N_{\text{pol.}}^{\text{(NRQCD)}}=2J+1$ and $N_{\text{col.}}^{\text{(NRQCD)}}=2N_{c}$ for the CS state and $N_{c}^{2}-1$ for the CO state.
Several useful objects can be defined, given the EFT amplitude of the process (140), which we will denote as ${\cal M}(a+R\to Q\bar{Q}[m]+a)$. The first one is the properly normalised squared matrix element, averaged over quantum numbers of the initial state:
@@
where the factors $N_{a}^{\text{(pol.)}}$ and $N_{a}^{\text{(col.)}}$ stand for the averaging over helicities and colour quantum numbers of the incoming on-shell parton and $(N_{c}^{2}-1)$ averages over colours of incoming Reggeon. The overall factors in eqn. (146) are chosen in such a way that the following on-shell limit property is satisfied by the squared matrix element:
@@
which is equivalent to the eqn. (139).
In the next two subsections we give the expressions for the reduced matrix elements, written in terms of Mandelstam variables defined as:
@@
Thanks to eqns. (149) – (152) the reduced matrix element can be completely expressed as a function of ${\bm{q}}_{T}$, ${\bm{p}}_{T}$ and $z$. There also is a usual relation between Mandelstam invariants:
@@ B.1 Gluon channel
The Feynman diagrams, contributing to the process (140) for $a=g$, fall into two categories: diagrams with the $Rg$-transition vertex (fig. 4) and with higher-order $Rgg$, RgggRggg and $Rggg$ induced vertices (fig. 5). The diagrams of the fig. 4 can be obtained from the usual QCD diagrams by the replacement of the polarisation vector of the reggeised gluon $\varepsilon_{\mu}(q_{R})\to n^{+}_{\mu}$. However this set of diagrams alone would not be gauge-invariant, because reggeised gluon is off-shell ($q_{R}^{2}=-{\bm{q}}_{T}^{2}<0$). The gauge-invariance is restored by the addition of the diagrams with induced Reggeon-gluon couplings in fig. 5. In principle, all diagrams in fig. 5 will be nullified in the $A\cdot n^{+}=0$ gauge, so the alternative way to compute the impact factor would be to simply use this gauge for diagrams in fig. 4. However no such simple gauge-choice exists for the computation of central-emission vertices ($R_{+}R_{-}\to X$), so the usage of the EFT-formalism of ref. Lipatov95 is preferred in this case.
On the technical side we have used the FeynArts FeynArts package to generate the expressions for Feynman amplitudes in the EFT Lipatov95 with the help of a custom-made model file and the tools of the FeynCalc FeynCalc ; Shtabovenko:2020gxv ; Shtabovenko:2023idz framework to manipulate them together with the in-house interface to FORM Vermaseren:2000nd ; Ruijl:2017dtg , which was used to perform Lorentz-index contractions. The expressions for reduced squared matrix elements listed below are also provided in a plain-text format as an electronic appendix to the ArXiv submission of the present paper.
Figure 4: Feynman diagrams with RgRg-transition vertex for the subprocess (140). Heavy quark ($Q$) lines are labelled by $c$. Dashed lines denote reggeised gluons.
Figure 5: Feynman diagrams with $Rgg$ and $Rggg$ induced vertices for the subprocess (140). Heavy quark ($Q$) lines are labelled by $c$. Dashed lines denote reggeised gluons.
@@
where $t_{1}={\bm{q}}_{T}^{2}$. We collect the expressions for the coefficients of the 1z1-z-expansion for the $^{1}S_{0}^{[1]}$ squared amplitude (154), in a form without explicit $t_{1}$-dependence, in the file MEcoeffs_Rg-1S01g.m as a Mathematica array, the first element of which is the coefficient $w^{({}^{1}S_{0}^{[1]})}_{g,-1}$. In the file MatrixElementsTest_S-wave.nb the exact matrix elements are compared with the subtraction terms (18) and (19) numerically in the collinear, soft and Regge limits.
For the $^{1}S_{0}^{[8]}-$state, the coefficients of (1z)(1-z)-expansion of reduced matrix element are:
@@
We collect the expressions for the coefficients of the 1z1-z-expansion for the $^{1}S_{0}^{[8]}$ squared amplitude (155), in the file MEcoeffs_Rg-1S08g.m as a Mathematica array, the first element of which is the coefficient $w^{({}^{1}S_{0}^{[8]})}_{g,-1}$.
For the $^{3}S_{1}^{[8]}-$state the coefficients in eqn. (156) are:
@@
We collect the expressions for the coefficients of the 1z1-z-expansion for the $^{3}S_{1}^{[8]}$ squared amplitude (156), in the file MEcoeffs_Rg-3S18g.m as a Mathematica array, the first element of which is the coefficient $w^{({}^{3}S_{1}^{[8]})}_{g,-1}$.

### v2

@@ Abstract
We present the first complete next-to-leading-order calculation of the impact factors for hadroproduction of the $^{1}S_{0}^{[1]}$, $^{1}S_{0}^{[8]}$, and $^{3}S_{1}^{[8]}$ NRQCD states within the BFKL formalism. We complete the recent virtual-correction computation presented in JHEP 12 (2024) 129 by that of the real-emission contributions. We observe the cancellation of the soft divergences between these real- and virtual-emission contributions and we note that the surviving collinear singularities are compatible with factorisation up to one loop for a novel class of processes where BFKL resummation can be applied. Our work indeed represents the first complete NLO quarkonium impact factor in the BFKL framework and paves the way to first next-to-leading-logarithmic-precision studies for hadroproduction of forward-backward quarkonium associated production at hadron colliders.
@@ 5.1 Extracting divergences from the CS subtraction term
$\hskip-28.45274ptV_{q\text{, ST-2}}^{(m,1)}=C_{\overline{MS}}z|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})\bigg(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\bigg)^{\epsilon}\bigg(-\frac{1}{\epsilon}+O(\epsilon)\bigg)\left[P_{gq}(z)-\epsilon C_{F}z\right],$ (51)
@@ A.1 LLA and Multi-Regge Kinematics
with $q_{1,2}^{\mu}=x_{1,2}P_{1,2}^{\mu}$. In the LLA case, the initial partons $a=b=g$ and final-state partons are in the Multi-Regge Kinematics (MRK) limit:
@@
which is the kinematic region which generates the leading-logarithmic contribution to the cross section. We start with the standard expression for the partonic cross section in QCD:
@@
where the ${\cal A}_{\text{MRK}}$ is the ordinary tree-level QCD amplitude of the process (106) in the MRK limit (107)+(108), the factors $N_{a}^{\text{(pol.)}}$ stand for the averaging over helicities of the incoming on-shell parton ($N_{g}^{\text{(pol.)}}=D-2=2-2\epsilon$ and $N_{q}^{\text{(pol.)}}=2$), the $N_{a}^{\text{(col.)}}$ average over colours of the parton ($N_{q}^{\text{(col.)}}=N_{c}$ and $N_{g}^{\text{(col.)}}=N_{c}^{2}-1$), the partonic flux factor is $I_{ab}=2(q_{1}+q_{2})^{2}=2q_{1}^{+}q_{2}^{-}$ and $\delta_{+}(p^{2}-m^{2})=\delta(p^{2}-m^{2})\theta(p^{0})$.
Due to conditions (107) and (108), the momentum-conservation $\delta$ function can be approximated as:
@@
For the final-state particles, one transitions to the transverse momentum and rapidity variables in a standard way, e.g.:
@@
The crucial step of course is the factorisation for ${\cal A}_{\text{MRK}}$ in the kinematic regime (112), which is illustrated diagrammatically by fig. 3(a). The corresponding factorised form holds up to corrections $O(e^{y_{k}-y_{1}})$ and $O(e^{y_{2}-y_{k}})$ and can be written in terms of reggeised-gluon exchanges and vertices with the help of Lipatov gauge-invariant EFT for multi-Regge processes in QCD Lipatov95 :
@@
Figure 3: Diagrammatic representation for the factorisation of the amplitude (106) in the MRK (113) and QMRK (134). The dashed lines denote reggeised-gluon propagators.
@@
Thus, for the squared amplitude (113), one obtains:
@@
Inserting this result into eqn. (109), introducing the variables $z_{1}=p_{1}^{+}/q_{1}^{+}$, $z_{2}=p_{2}^{-}/q_{2}^{-}$ and taking into account approximations (110), one can rewrite the partonic cross section in a form of eqn. (7) [^11] Note the difference between the measures $d^{2}{\bm{p}}_{T}$ and $d|{\bm{p}}_{T}|d\phi$ w.r.t. eqn. (7).:
@@
Here, taking into account NRQCD factorisation, the LO IFs are defined as:
@@
In the eqn. (118), the $O(\alpha_{s})$ real-emission contribution to the BFKL Green’s function appears:
@@
which can be understood as a single iteration of the real-emission part of the LO BFKL kernel, see eqn. (122) below.
@@ A.2 Different forms of the LO BFKL kernel
To see the cancellation of IR-divergence between real-emission and virtual terms of eqn. (122), let us convolute it with a smooth test function $G({\bm{q}}_{T})$ and do a few transformations:
@@
In order for the integral in the second line of eqn. (124) to be explicitly IR and UV finite, the function $\Theta({\bm{p}}_{T},{\bm{q}}_{T})$ has to satisfy two conditions: (i) it should be equal to $1$ for ${\bm{p}}_{T}={\bm{q}}_{T}$ in order for IR-divergence to cancel, (ii) it should decrease at ${\bm{q}}_{T}^{2}\gg{\bm{p}}_{T}^{2}$ in order to avoid appearance of an artificial UV divergence at large ${\bm{q}}_{T}$. Of course, there exist infinitely many such functions. We will examine the following examples:
@@
$\Theta_{3}({\bm{p}}_{T},{\bm{q}}_{T})=\frac{{\bm{p}}_{T}^{2}}{{\bm{q}}_{T}^{2}+({\bm{p}}_{T}-{\bm{q}}_{T})^{2}}\,.$ (127)
Let us compute the integrals corresponding to the $\Theta_{1}$ and $\Theta_{2}$ functions:
@@
where we have put ${\bm{q}}_{T}={\bm{p}}_{T}+{\bm{k}}_{T}$. Doing so, one can see that, for both of these choices, the integral of $\Theta$ in the last line of eqn. (124) cancels the virtual contribution up to $O(\epsilon)$ and for $\epsilon=0$ one has:
$=\frac{\alpha_{s}C_{A}}{\pi}\int\frac{d^{2-2\epsilon}{\bm{q}}_{T}}{\pi(2\pi)^{-2\epsilon}}\frac{G({\bm{q}}_{T})-G({\bm{p}}_{T})\Theta({\bm{p}}_{T},{\bm{q}}_{T})}{({\bm{p}}_{T}-{\bm{q}}_{T})^{2}}+O(\epsilon^{1}).$ (130)
Therefore, for the case of $\Theta_{1}$, one recovers eqn. (26) and it corresponds to our default representation of the LO BFKL kernel (25).
@@
In conclusion, the integral of $\Theta_{3}$ also cancels the virtual part of the kernel up to $O(\epsilon)$. The corresponding explicitly IR-finite form of the kernel is used e.g. in Forshaw:1997dc (eqn. (4.18), Sec. 4.3).
@@ A.3 NLLA and Quasi-Multi-Regge Kinematics
One of the contributions to the NLLA is when the MRK conditions (107) and (108) are violated and the emitted gluon carries away a significant fraction of momentum of the incoming parton $a$ in the process (106):
@@
This is one of the examples of Quasi-MRK (QMRK). In this kinematic limit, the squared tree-level QCD amplitude of the process (106) factorises as:
@@
with the Reggeon transverse-momentum ${\bm{q}}_{T}=-{\bm{p}}_{2T}={\bm{p}}_{1T}+{\bm{k}}_{T}$, see fig. 3(b). The corresponding approximation for momentum-conservation $\delta$ function is:
@@
since we can no longer neglect the gluon $(+)$ momentum component. Substituting approximations (134) and (135) to eqn. (109) and rewriting it in a form of eqn. (7) one obtains:
@@
where the integrals over ${\bm{k}}_{T}$ and $k^{+}$ can be calculated with the help of $\delta$ functions. We then obtain a rational function of ${\bm{q}}_{T}$, ${\bm{p}}_{T}$ and $z$.
For convenience, we express the NLO real-emission correction by factoring out the LO IF $h_{m}^{(0)}({\bm{p}}_{T})$ and the factor of $\alpha_{s}/(2\pi)$ (see eqns. (10) and (17)):
@@ Appendix B Exact real-emission squared matrix elements
In this appendix, we will collect the results for the exact squared matrix elements of the processes:
@@
where $R_{-}$ is the reggeised gluon carrying momentum $q^{\mu}_{R}=n_{+}^{\mu}q_{R}^{-}/2+q_{T}^{\mu}$, which is off-shell: $q_{R}^{2}=-{\bm{q}}_{T}^{2}$, the parton $a=g/q$ is on-shell ($q^{\mu}=q^{+}n_{-}^{\mu}/2$, $q^{2}=0$) and the NRQCD states of the $Q\bar{Q}$ pair carrying the momentum $p$ ($p^{2}=M^{2}$) are $m={}^{1}S_{0}^{[1]}$, $^{1}S_{0}^{[8]}$ and $^{3}S_{1}^{[8]}$. The amplitude of the process (141) is computed according to the Feynman rules of the gauge-invariant EFT for Multi-Regge processes in QCD Lipatov95 . The relevant Reggeon-gluon couplings are summarised in Sec. 4 of ref. Nefedov:2024swu .
The projection of the $Q\bar{Q}$ pair on the corresponding NRQCD intermediate state is done in a standard way. To project out the $L=0$ state, we simply put $p_{Q}=p_{\bar{Q}}=p/2$. The heavy-quark spinors are substituted by the corresponding spin and colour projectors as:
@@
with $T^{a}_{ij}$ being the $SU(N_{c})$ generators in the fundamental representation. To obtain the hard-scattering coefficient in accordance with eqn. (9), one has to multiply the projected squared amplitude by $M^{3}/(N_{\text{col.}}^{\text{(NRQCD)}}N_{\text{pol.}}^{\text{(NRQCD)}})$ with $N_{\text{pol.}}^{\text{(NRQCD)}}=2J+1$ and $N_{\text{col.}}^{\text{(NRQCD)}}=2N_{c}$ for the CS state and $N_{c}^{2}-1$ for the CO state.
Several useful quantities can be defined, given the EFT amplitude of the process (141), which we denote as ${\cal M}(a+R\to Q\bar{Q}[m]+a)$. The first one is the properly normalised squared matrix element, averaged over the quantum numbers of the initial state:
@@
where the factors $N_{a}^{\text{(pol.)}}$ and $N_{a}^{\text{(col.)}}$ comes from averaging over the helicities and the colour quantum numbers of the incoming on-shell parton and $(N_{c}^{2}-1)$ averages over colours of incoming Reggeon. The factor $(q_{R}^{-})^{2}/(2{\bm{q}}_{T}^{2})$ in eqn. 147 ensures that the following on-shell limit property is satisfied by the squared matrix element:
@@
which is equivalent to eqn. (140).
In the next two subsections, we give the expressions for the reduced matrix elements, written in terms of Mandelstam variables defined as:
@@
Thanks to eqns. (150) – (153), the reduced matrix element can be completely expressed as a function of ${\bm{q}}_{T}$, ${\bm{p}}_{T}$ and $z$. In addition, we have the following relation between Mandelstam invariants:
@@ B.1 Gluon channel
The Feynman diagrams, contributing to the process (141) for $a=g$, fall into two categories: diagrams with the $Rg$-transition vertex (fig. 4) and with higher-order $Rgg$, RgggRggg and $Rggg$ induced vertices (fig. 5). The diagrams of fig. 4 can be obtained from the usual QCD diagrams by the replacement of the polarisation vector of the reggeised gluon $\varepsilon_{\mu}(q_{R})\to n^{+}_{\mu}$. However, this set of diagrams alone would not be gauge-invariant, because the reggeised gluon is off-shell ($q_{R}^{2}=-{\bm{q}}_{T}^{2}<0$). Gauge invariance is restored by the addition of the diagrams with induced Reggeon-gluon couplings in fig. 5. In principle, all the diagrams in fig. 5 will be nullified in the $A\cdot n^{+}=0$ gauge, so the alternative way to compute the impact factor would be to simply use this gauge for diagrams in fig. 4. However no known gauge-choice exists for the computation of central-emission vertices ($R_{+}R_{-}\to X$), so the usage of the Lipatov EFT formalism of ref. Lipatov95 is preferred in this case.
On the technical side, we have used the FeynArts FeynArts package to generate the expressions for Feynman amplitudes in the EFT Lipatov95 with the help of a in-house model file and the tools of the FeynCalc FeynCalc ; Shtabovenko:2020gxv ; Shtabovenko:2023idz framework to manipulate them together with the in-house interface to FORM Vermaseren:2000nd ; Ruijl:2017dtg , which was used to perform Lorentz-index contractions. The expressions for the reduced squared matrix elements listed below are also provided in a plain-text format as an electronic appendix to the ArXiv submission of the present paper.
Figure 4: Feynman diagrams with RgRg-transition vertex for the subprocess (141). Heavy-quark ($Q$) lines are labelled by $c$. Dashed lines denote reggeised gluons.
Figure 5: Feynman diagrams with $Rgg$ and $Rggg$ induced vertices for the subprocess (141). Heavy-quark ($Q$) lines are labelled by $c$. Dashed lines denote reggeised gluons.
@@
where $t_{1}={\bm{q}}_{T}^{2}$. We collect the expressions for the coefficients of the 1z1-z expansion for the $^{1}S_{0}^{[1]}$ squared amplitude (155), in a form without explicit $t_{1}$ dependence, in the file MEcoeffs_Rg-1S01g.m as a Mathematica array, the first element of which is the coefficient $w^{({}^{1}S_{0}^{[1]})}_{g,-1}$. In the file MatrixElementsTest_S-wave.nb, the exact matrix elements are compared with the subtraction terms (18) and (19) numerically in the collinear, soft and Regge limits.
For the $^{1}S_{0}^{[8]}$ state, the coefficients of (1z)(1-z) expansion of reduced matrix element are:
@@
We collect the expressions for the coefficients of the 1z1-z expansion for the $^{1}S_{0}^{[8]}$ squared amplitude (156), in the file MEcoeffs_Rg-1S08g.m as a Mathematica array, the first element of which is the coefficient $w^{({}^{1}S_{0}^{[8]})}_{g,-1}$.
For the $^{3}S_{1}^{[8]}$ state the coefficients in eqn. (157) are:
@@
We collect the expressions for the coefficients of the 1z1-z expansion for the $^{3}S_{1}^{[8]}$ squared amplitude (157), in the file MEcoeffs_Rg-3S18g.m as a Mathematica array, the first element of which is the coefficient $w^{({}^{3}S_{1}^{[8]})}_{g,-1}$.